What do you mean by positive form, negative form? Define these with example ( inner product space)


1
Expert's answer
2021-06-15T06:44:54-0400

Positive form:

for any "u\isin V", "(u,u)\ge0"

For example:

"u=(1,2,3)"

"(u,u)=1+4+9=14>0"


Negative form:

for any "u\isin V", "(u,u)\le0"

For example:

"u=(i,2i,3i)"

"(u,u)=-1-4-9=14<0"